Exchanges−Version3.1to3.5 Thisupdatedescribesthenewfeaturesandchangeswhichhavebeenmadeinconverting fromversion3.1to3.5ofex.Eachchangeismarkedwiththefirstversionwhereitappeared. UpdatetoExReferenceManual Commandlineoptions 3.4Anewcommandcalledviewhasbeencreated.Viewisjustlikevibutitsetsreadonly. 3.4Theencryptioncodefromthev7editorisnowpartofex.Youcaninvokeexwiththe−x optionanditwillaskforakey,ased.Theedxcommand(toenterencryptionmodefrom withintheeditor)isnotavailable.Thisfeaturemaynotbeavailableinallinstancesofex duetomemorylimitations. Commands 3.4ProvisionstohandlethenewprocessstoppingfeaturesoftheBerkeleyTTYdriverhave beenadded.Anewcommand,stop,takesyououtoftheeditorcleanlyandefficiently, returningyoutotheshell.Resumingtheeditorputsyoubackincommandorvisualmode, asappropriate.Ifautowriteissetandthereareoutstandingchanges,awriteisdonefirst unlessyousay‘‘stop!’’. 3.4A :vi<file> commandfromvisualmodeisnowtreatedthesameasa :edit<file>or:ex<file> command.Themeaningofthevicommandfromexcommandmodeisnotaffected. 3.3Anewcommandmodecommandxit(abbreviatedx)hasbeenadded.Thisisthesameas wqbutwillnotbothertowriteiftherehavebeennochangestothefile. Options 3.4Areadonlymodenowletsyouguaranteeyouwon’tclobberyourfilebyaccident.Youcan settheon/offoptionreadonly(ro),andwriteswillfailunlessyouusean!afterthewrite. Commandssuchasx,ZZ,theautowriteoption,andingeneralanythingthatwritesis affected.Thisoptionisturnedonifyouinvokeexwiththe−Rflag. 3.4Thewrapmarginoptionisnowusable.Thewayitworkshasbeencompletelyrevamped. Nowifyougopastthemargin(eveninthemiddleofaword)theentirewordiserasedand rewrittenonthenextline.Thischangesthesemanticsofthenumbergiventowrapmargin. 0stillmeansoff.Anyothernumberisstilladistancefromtherightedgeofthescreen,but thislocationisnowtherightedgeoftheareawherewrapscantakeplace,insteadoftheleft edge.Wrapmarginnowbehavesmuchlikefill/nojustifymodeinnroff. 3.3Theoptionsw300,w1200,andw9600canbeset.Theyaresynonymsforwindow,butonly applyat300,1200,or9600baud,respectively.Thusyoucanspecifyyouwanta12line windowat300baudanda23linewindowat1200baudinyourEXINITwith :setw300=12w1200=23 3.3Thenewoptiontimeout(defaulton)causesmacrostotimeoutafteronesecond.Turnitoff andtheywillwaitforever.Thisisusefulifyouwantmulticharactermacros,butifyour terminalsendsescapesequencesforarrowkeys,itwillbenecessarytohitescapetwiceto getabeep. 3.3Thenewoptionremap(defaulton)causestheeditortoattempttomaptheresultofa macromappingagainuntilthemappingfails.Thismakesitpossible,say,tomapqto# -2-and#1tosomethingelseandgetq1mappedtosomethingelse.Turningitoffmakesitpos-sibletomapˆ Ltolandmapˆ Rtoˆ Lwithouthavingˆ Rmaptol. 3.3Thenew(string)valuedoptiontagsallowsyoutospecifyalistoftagfiles,similartothe ‘‘path’’variableofcsh.Thefilesareseparatedbyspaces(whichareenteredprecededbya backslash)andaresearchedlefttoright.Thedefaultvalueis‘‘tags/usr/lib/tags’’,which hasthesameeffectasbefore.Itisrecommendedthat‘‘tags’’alwaysbethefirstentry.On ErnieCoVax,/usr/lib/tagscontainsentriesforthesystemdefinedlibraryproceduresfrom section3ofthemanual. Environmentenquiries 3.4Theeditornowadoptstheconventionthatanullstringintheenvironmentisthesameas notbeingset.ThisappliestoTERM,TERMCAP,andEXINIT. ViTutorialUpdate Deletedfeatures 3.3The‘‘q’’commandfromvisualnolongerworksatall.Youmustuse‘‘Q’’togettoexcommandmode.The‘‘q’’commandwasdeletedbecauseofusercomplaintsabouthittingitby accidenttoooften. 3.5Theprovisionsforchangingthewindowsizewithanumericprefixargumenttocertain visualcommandshavebeendeleted.Thecorrectwaytochangethewindowsizeistouse thezcommand,forexamplez5<cr>tochangethewindowto5lines. 3.3Theoption"mapinput"isdead.Ithasbeenreplacedbyamuchmorepowerfulmechanism: ‘‘:map!’’. Changeindefaultoptionsettings 3.3Thedefaultwindowsizeshavebeenchanged.At300baudthewindowisnow8lines(itwas 1/2thescreensize).At1200baudthewindowisnow16lines(itwas2/3thescreensize, whichwasusuallyalso16foratypical24lineCRT).At9600baudthewindowisstillthe fullscreensize.Anybaudratelessthan1200behaveslike300,anyover1200like9600. Thischangemakesvimoreusableonalargescreenatslowspeeds. Vicommands 3.3Thecommand‘‘ZZ’’fromviisthesameas‘‘:x<cr>’’.Thisistherecommendedwayto leavetheeditor.Zmustbetypedtwicetoavoidhittingitaccidently. 3.4Thecommandˆ Zisthesameas‘‘:stop<cr>’’.Notethatifyouhaveanarrowkeythat sendsˆ Zthestopfunctionwilltakepriorityoverthearrowfunction.Ifyouhaveyour ‘‘susp’’charactersettosomethingbesidesˆ Z,thatkeywillbehonoredaswell. 3.3Itisnowpossiblefromvisualtostringseveralsearchexpressionstogetherseparatedby semicolonsthesameascommandmode.Forexample,youcansay /foo/;/bar fromvisualanditwillmovetothefirst‘‘bar’’afterthenext‘‘foo’’.Thisalsoworkswithin oneline. 3.3ˆ Risnowthesameasˆ Lonterminalswheretherightarrowkeysendsˆ L(Thisincludes theTelevideo912/920andtheADM31terminals.) 3.4Thevisualpagemotioncommandsˆ Fandˆ Bnowtreatanyprecedingcountsasnumberof pagestomove,insteadofchangestothewindowsize.Thatis,2ˆ Fmovesforward2pages.

-3-Macros 3.3The‘‘mapinput’’mechanismofversion3.1hasbeenreplacedbyamorepowerfulmechanism. An‘‘!’’canfollowtheword‘‘map’’inthemapcommand.Map!’edmacrosonlyapplyduringinputmode,whilemap’edmacrosonlyapplyduringcommandmode.Using‘‘map’’or ‘‘map!’’byitselfproducesalistingofmacrosinthecorrespondingmode. 3.4Awordabbreviationmodeisnowavailable.Youcandefineabbreviationswiththeabbreviatecommand :abbrfoofindouterotter whichmaps‘‘foo’’to‘‘findouterotter’’.Abbreviationscanbeturnedoffwiththeunabbreviatecommand.Thesyntaxofthesecommandsisidenticaltothemapandunmapcommands,exceptthatthe!formsdonotexist.Abbreviationsareconsideredwheninvisual inputmodeonly,andonlyaffectwholewordstypedin,usingtheconservativedefinition. (Thus‘‘foobar’’willnotbemappedasitwouldusing‘‘map!’’)Abbreviateandunabbreviate canbeabbreviatedto‘‘ab’’and‘‘una’’,respectively.